OpenStudy (anonymous):

It takes 6 hours for a plane to travel 720 km with a tail wind and 8 hours to make the return trip with a head wind. Find the airspeed of the plane and speed of the wind current

OpenStudy (anonymous):

What are tail/head winds?

hero (hero):

We know this is a distance, rate, time problem so d = rt is the appropriate formula to use. However since we have to consider both airspeed and wind speed, we'll have to alter the formula a bit. a = airspeed w = windspeed \(r = a \pm w\)

hero (hero):

So we'll use the formula \(d = (a \pm w)t\) In other words, we need to use two separate formulas one for headwind, and the other for tailwind: \(d = (a - w)t\) \(d = (a + w)t\)

hero (hero):

d = 720 For headwind, t = 8 For tailwind, t = 6 so: \(720 = (a - w)8\) \(720 = (a + w)6\)

hero (hero):

Now divide the first equation by 8 Divide the second equation by 6 \(90 = a - w\) \(120 = a + w\)

hero (hero):

Now you have a systems of equations. You can finish solving using elimination method. Try solving for a

OpenStudy (anonymous):

THANK YOU!!!!!!!!!!!!!!! YOU ARE MY FAVORITE PERSON EVER!!!!!

hero (hero):

What values did you get for a and w?

OpenStudy (anonymous):

a=105 and w=15?

hero (hero):

Good job